OpenStudy (anonymous):

*fan and medal awarded* The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today? 900 620 520 450

OpenStudy (anonymous):

This answer will only tell you the number of one-step equations that have been eliminated. We don't actually know how many auditioned. Also, be careful with the use of ! in mathematics as 1070! = 1070 x 1069 x 1068 x.......x 3 x 2 x1 Let A be the number of one-step equations eliminated and B the number of two-step equations eliminated then, A + B = 1070 .....Equation (1) and we are told that, 3A - 2B = 110.....Equation (2) multiplying Eqn (1) by 2 gives 2A + 2B = 2140........now add Eqn(2) 3A - 2B = 110 5A = 2250 ⇒ A = 450....which is the number of one-step equations ELIMINATED today.

OpenStudy (anonymous):

thanks sooo much !!!!!!

OpenStudy (anonymous):

would u mind helping me with another ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

bring it on